Seven people go to a cinema and park their car outside,which are identical.while returning only three of them were able to unlock their cars.find total no. of ways in which this can be done
 

Grade:12

2 Answers

Vijay Mukati
8 years ago
This will be equal to selecting 3 out of 7 people and then rearrange them but subtracting one case for which the cars will be unlock. 7P3 -1.
mahein
2 years ago
choose 3 out of 7 and dearrange rest 4 because only three of them are right so answer will be 7C3*D9=7*5*9=315